OpenStudy (anonymous):

Mike bought an MP3 player for $150. The sales-tax rate was 8 percent. How much sales tax did he pay on the MP3 player?

OpenStudy (anonymous):

no i suck at math.. :/

OpenStudy (anonymous):

Wait its 12 right?

OpenStudy (anonymous):

I'm thinking they are looking for a different answer.

OpenStudy (anonymous):

help me on this one please :3 Jan bought a blazer for $39.90. The price of the blazer was $38. What sales-tax rate was she charged for this blazer?

OpenStudy (anonymous):

$150 is the total price including tax So the total cost = price + tax Let P = Price And P (.08) = tax So equation is P + .08P = 150 P (1+.08) =150 P = 138.89 So the tax = 150-138.89 or 11.11

OpenStudy (anonymous):

$158 dollars. Just fins 8 percent of 150 dollars and then subtract that amount from 150 dollars

OpenStudy (anonymous):

On the 2nd one, it's easier. Price is $38, tax is $1.50. Tax Rate = tax/price = 1.50/38 =3.95%

OpenStudy (anonymous):

Make sure you look at my longer solution for the 1st one. The tax is not on the total cost. It's on the price of the item and what you pay for something is price + tax.

OpenStudy (anonymous):

i know i got it thanks, :) do you mind helping me on this one.. Rita bought a pair of jeans for $89. The sales-tax rate is 11 percent. What is the total amount she paid for the jeans?

OpenStudy (anonymous):

79.21 dollars

OpenStudy (anonymous):

So, this is the 3rd way they can ask this question. They are a little confusing, you have to read them carefully to see what they are asking. Here it sounds like, the jeans cost $89 and you have to add the tax to it, and see what the TOTAL AMOUNT is. So, cost = $89 tax = $89 * .11 = $9.79 TOTAL AMOUNT = $98.79 You now have the 3 different ways this question can be asked.

OpenStudy (anonymous):

so wait its like all the same just written differently?

OpenStudy (anonymous):

No when you find the tax price you subtract that from the other price which is 89 dollars you don't add

OpenStudy (anonymous):

@Tanzina....they are asking for the TOTAL AMOUNT....so the one given can't be the total...it has to be the price of the item BEFORE tax.

OpenStudy (anonymous):

Oh ya at first I didn't know that your right sorry How do I give you a medal

OpenStudy (anonymous):

@ceceee In each question, they gave different information for the same general problem. Price of item * tax = Total Amount in #1) they gave tax rate and total cost and you had to solve for tax amount in #2) they gave price of item and total cost, and you had to solve for tax rate in #3) they gave price of item and tax rate and you had to solve for total amount You have to read the question carefully and figure what they have given you and what they are asking for.

OpenStudy (anonymous):

@Tanzina no worries, the wording on these types of questions is usually poor and very easy to get wrong if you read it quickly

OpenStudy (anonymous):

You see aim in advanced regents placement class and I forgot how to do this

OpenStudy (anonymous):

@Scott* okay so is this the same also? Kate is at a T-shirt sale where she can buy one and get a second one for 30% off. She wants to buy two T-shirts listed at $19 each. How much will she pay?

OpenStudy (anonymous):

Well this is a different question. You have to read carefully what the question is asking. This one says T-shirts cost $19 each, but the 2nd one is 30% off. So, T-shirt #1 costs $19 T-shirt #2 costs $19 less 30% of $19 or $13.30 $19 + $13.30 = $32.3

OpenStudy (anonymous):

okay im still confuse and im really sorry if i seem dumb but just i need to pass this math test to go back to my old school. ..so can you please help me on this one Mary is buying several items that cost $128.25 total. She is using a store coupon for 35% off her purchases. She has to pay 4% sales tax. Calculate the total cost of the items.

OpenStudy (anonymous):

I hate the way they use the word cost to mean multiple things. It is confusing.

OpenStudy (anonymous):

Let's step thru this one....it's actually straight forward. So, what is the price tag (before tax and discount) of everything she bought?

OpenStudy (anonymous):

19, but for both without the discount is 38 ... correct?

OpenStudy (anonymous):

wait...just from the question....read it again....no math...what is the price tag?

OpenStudy (anonymous):

ohgosh sorry >.< i thought it wasanother question sorry. but 128.25?

OpenStudy (anonymous):

Right...now do you apply the discount first or sales tax?

OpenStudy (anonymous):

the sales tax?

OpenStudy (anonymous):

No...think about when you go into a store...an item is $20, but it's on sale for $10....do you pay tax on the $20 or $10.

OpenStudy (anonymous):

the 10

OpenStudy (anonymous):

Correct, after the discount...so, let's calculate what the discount is and what the new price of the items are....go ahead and try

OpenStudy (anonymous):

is the new price 83.36?

OpenStudy (anonymous):

Yes...nice job...now what do you have to do?

OpenStudy (anonymous):

i have a lot to do ;c

OpenStudy (anonymous):

sorry...I meant with this question...you haven't finished the question yet

OpenStudy (anonymous):

oh lol sorry, hm you find out the sales tax?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and is that 4.00?

OpenStudy (anonymous):

No...it's 4% of the price of the items

OpenStudy (anonymous):

hm idk how to do that ..

OpenStudy (anonymous):

wait is it 86.69?

OpenStudy (anonymous):

Price of the items * tax rate = tax amount So, 83.36 * .04 = 3.33 Now, the total cost is 83.36 + 3.33 = 86.69

OpenStudy (anonymous):

Nice job

OpenStudy (anonymous):

thank you :) is calculating tips hard?

OpenStudy (anonymous):

You know how to do the math...now you just have to read very carefully the word problem and figure out what they are asking you

OpenStudy (anonymous):

No problem..... Calculating tips is same type of math. Let's use easy numbers: You spend $50 on dinner and want to give a 10% tip, You give $5. Now, the hardest question they can ask is: You spend a total of $55 on dinner and tipped 10%, how much did you tip?

OpenStudy (anonymous):

you tip 10 dollars?

OpenStudy (anonymous):

Look at the first one...the answer is $5... now how do you get there? Do you know the distributive property: A(B +C) = AB + AC ???

OpenStudy (anonymous):

You are confusing % and dollars...If they say % you have to multiply that by something to get how many $$

OpenStudy (anonymous):

okay i get it now. do you just subtract half from the percentage? and yes i do

OpenStudy (anonymous):

OK...this question isn't straight forward...it's like the 1st one you asked in this thread....I'll explain

OpenStudy (anonymous):

Let P = Price of meal Let T% = percent you want to tip Let T = $$ you will tip Let C = Total Cost of the meal P + T = C (price of meal + $ you will tip = total cost T = T% * P ($ you will tip = price of meal * tip percentage so P + (T% * P) = C using distributive property you can take out the P and get: P (1 + T%) = C So, in the question I asked: C = $55 T% = 10% P(1+10%)=$55 1.1P = 55 (divide both sides by 1.1) P = 50 you can calculate tip 2 ways: Total Cost - Price of meal = 55 - 50 = 5 or Tax % * price of meal = 10% * 50 = 5 There is a lot there. But if you can handle this one, you will be able to handle them all.

OpenStudy (anonymous):

THANK YOU ! i saved this, you broke it down easier

OpenStudy (anonymous):

Good luck...I'm taking off...if you need anything continue posting...either someone else will pickup and or I'll check back in tomorrow for a little bit.

OpenStudy (anonymous):

okay thank you again :)